OpenStudy (anonymous):

Help?! Two trains leave the station at the same time, one heading west and the other east. The westbound train travels 16 miles per hour slower than the eastbound train. If the two trains are 680 miles apart after 4 hours, what is the rate of the westbound train? Do not do any rounding.

OpenStudy (anonymous):

@micahm

OpenStudy (micahm):

name of lesson

OpenStudy (anonymous):

its like speed/distance or something i dont know i dont pay much attention to it

OpenStudy (micahm):

have you seen family ties cus this question was on it with Michael j fox 360=(X+X-14)*2 360=(2X-14)2 360=4X-28 4X=360+28 4X=388 X=388/4 X=97 mph. speed of the faster train. 97-14=83 mph. speed of the slower train. Proof: 360=(97+83)*2 360=180*2 360=360

OpenStudy (anonymous):

so the answer is 83?

OpenStudy (micahm):

most like-ly yes but look at again before you choose

OpenStudy (anonymous):

it wasnt right

OpenStudy (micahm):

that's why you look before you choose

OpenStudy (anonymous):

the answer was 77

OpenStudy (micahm):

this problem i paste was to help you solve it